OpenStudy (anonymous):

Will MEDAL Cos(sin2/3 + tan3/2) Sum and Difference Formula Can someone do this problem and tell me what they get so I can compare it with my own

OpenStudy (usukidoll):

so.. we're dealing with \[\cos(\theta+ \beta) \] where \[\theta = \sin \frac{2}{3}, \beta = \tan \frac{3}{2}\] ?

OpenStudy (usukidoll):

\[\cos( \theta+ \beta) = \cos(\theta)\cos(\beta)-\sin(\theta)\sin(\beta)\]

OpenStudy (usukidoll):



OpenStudy (usukidoll):

if my beta is tangent then I have to find cosine and sine of my beta